(a.) Explain the term Doppler effect.

  A car sounding an alarm of frequency 512Hz is approaching a stationary car at a speed of 10ms-1

 (b.) Explain why the driver of the stationary car has the impression that the frequency of the sound is

    increasing.



1
Expert's answer
2021-03-21T11:32:43-0400

(a) Doppler Effect refers to the change in wave frequency during the relative motion between a wave source and its observer.



"f_I=\dfrac{340}{340-10}\times 512=526.34\ Hz"


"\\ f_P=\dfrac{340+10}{340}\times 526.34=537.68\ Hz"


(b) Sound is a longitudinal pressure wave. As the alarms distance became less , the air molecules get compressed together. The wavelength of the sound (these pressure waves) decreases, and the frequency increases. That results in a higher sound pitch